OpenStudy (anonymous):

solve using the rule of exponents...(a^1/2 b ^-1/3)(ab)

jigglypuff314 (jigglypuff314):

when multiplying exponents, you add the exponents so a^1/2 * a = a^((1/2)+1) and b^(-1/3)*b = b^((-1/3)+1)

OpenStudy (jdoe0001):

\(\bf \large {(a^{\frac{1}{2}}b^{-\frac{1}{3}})(ab)\implies(a^{\frac{1}{2}}b^{-\frac{1}{3}})(a^1b^1)\implies a^{\frac{1}{2}}\cdot a^1\cdot b^{-\frac{1}{3}}\cdot b^1 \\ \quad \\ \textit{keep in mind that }\qquad x^n\cdot x^m\implies x^{n+m}}\)

OpenStudy (anonymous):

I"m still very confused